### v2

# Exploring Generative Process Reward Modeling for Semi-Structured Data: A Case Study of Table Question Answering

Abstract
Process reward models (PRMs) enhance complex reasoning in large language models (LLMs) by evaluating candidate solutions step-by-step and selecting answers based on aggregated step scores. While effective in domains such as mathematics, their applicability to tasks involving semi-structured data, like table question answering (TQA), remains unexplored. TQA poses unique challenges for PRMs, including abundant irrelevant information, loosely connected reasoning steps, and domain-specific reasoning. This work presents the first systematic study of PRMs for TQA. We evaluate state-of-the-art generative PRMs on TQA from both answer and step perspectives. Results show that PRMs that combine textual and code verification can aid solution selection but struggle to generalize to out-of-domain data. Analysis reveals a weak correlation between performance in step-level verification and answer accuracy, possibly stemming from weak step dependencies and loose causal links. Our findings highlight limitations of current PRMs on TQA and offer valuable insights for building more robust, process-aware verifiers.
Exploring Generative Process Reward Modeling for Semi-Structured Data: A Case Study of Table Question Answering
Lei Tang2 Wei Zhou1,3 Mohsen Mesgar3 1University of Augsburg, Germany 2Institut für Maschinelle Sprachverarbeitung, University of Stuttgart, Germany 3Bosch Center for Artificial Intelligence, Renningen, Germany letang9-c@my.cityu.edu.hk {wei.zhou3|mohsen.mesgar}@de.bosch.com
1 Introduction
Recent advances in process reward models (PRMs) have demonstrated remarkable improvements in the reasoning capabilities of large language models (LLMs) Wang et al. (2024b); Yin et al. (2025); Ong et al. (2025). PRMs enable sampling-based test time scaling (TTS) by grading multiple solutions step-wise. The best solution is then selected based on aggregated step scores, e.g., average step scores. This verification-driven paradigm has proven highly effective in domains such as math Zhao et al. (2025); Zhang et al. (2025b, a), visual Ong et al. (2025); Wang et al. (2025), and textual QA Yin et al. (2025). However, to the best of our knowledge, no previous work explores the effectiveness of PRMs on tasks using semi-structured data, e.g., tables. This raises a natural and important question: Is the success of PRMs in domains like math transferable to other complex reasoning tasks, where semi-structured data are involved? To answer the question, we focus on table question answering (TQA), where a system needs to solve a question based on information in a table.
TQA differs fundamentally from math problems. In math, most given information is relevant, and reasoning steps are tightly linked in deductive chains. In contrast, tables in TQA can be large with abundant irrelevant information for solving a question. Moreover, the reasoning steps are often loosely connected, i.e., later steps may not strictly follow from earlier ones and can start a new topic. Consider “Step 1” and “Step 2” on the right side of Figure 1. Nevertheless, both tasks require multi-step and complex reasoning to be solved.
We hypothesize that TQA can benefit from PRM-based verification. Unlike existing TQA approaches, which predominantly optimize policy models through prompt engineering Zhou et al. (2025); Wang et al. (2024c), or fine-tuning Zhang et al. (2024); Deng and Mihalcea (2025), PRMs are model agnostic: they can be applied to outputs from any policy models and rewards given by PRM can be used to further improve performance of policy model via reinforcement learning with verifiable rewards (RLVR) Su et al. (2025). Moreover, solutions selected by PRMs reflect both answer and step quality, compared to majority voting that considers only answer frequencies (Liu et al., 2024; Zhou et al., 2025). Note that the effectiveness of majority voting depends strongly on the base model’s capability: it has to produce the correct answer in most sampled outputs. By contrast, PRMs are specifically trained for step-level evaluation. Last but not least, PRMs can enhance system trustworthiness, providing human-readable rationales for verification decisions on step correctness.
Yet, applying PRMs to TQA is non-trivial. Retrieval noise, loosely connected reasoning steps, and domain-specific reasoning patterns may limit their effectiveness. As an initial step, this work assesses the potential of applying PRMs for step-level verification in TQA. We explore to what extent such a step-based evaluation paradigm assists in solution selection, with a focus on PRMs that generate human-understandable rationales. Specifically, we evaluate multiple PRM variants from both answer-level and step-level perspectives against strong baselines, including self-consistency Wang et al. (2022) and LLM-as-a-judge Gu et al. (2024). Our results show that PRMs can aid in selecting high-quality solutions for TQA, particularly when combining textual and code verification. However, their effectiveness remains limited, with notable weaknesses such as poor generalization to out-of-domain datasets.
Through qualitative and quantitative analysis of verification chains, we uncover a weak correlation between step verification and answer correctness: better average step scores do not necessarily yield higher answer accuracy in TQA. This stems from characteristics of the problem profile, such as weak step dependence and loose causal relationships between reasoning steps. By bridging PRMs with TQA, our work clarifies when and why PRMs succeed or fail, and provides insights for designing more robust, process-aware systems for complex multi-step reasoning tasks. Code is available.111https://github.com/boschresearch/TQA-PRM
2 Examining PRMs on TQA.
This section describes the PRMs we examine, the training setup, baselines, TQA datasets for training and evaluation, as well as evaluation protocols.
Problem Formulation.
Let denote a table, a natural language question, and a set of candidate reasoning paths generated by a policy model , where each path produces a predicted answer . A PRM evaluates each reasoning step by generating a natural language rationale and assigning a scalar reward . The goal is to identify the reasoning path that maximizes the likelihood of producing a correct answer, i.e., , where is the ground-truth answer.
Examined PRMs.
We focus on state-of-the-art generative PRMs that output both step-level rewards and rationales explaining these rewards. GenPRM Zhao et al. (2025) verifies each step through a two-stage process: (1) a textual verification stage, in which the PRM examines a step using chain-of-thought (CoT) reasoning Wei et al. (2023), and (2) a code verification stage, where Python code is generated and executed for the current step. Step correctness is determined based on the CoT reasoning, generated code, and its execution results. Textual PRM Zhao et al. (2025) is an ablation of GenPRM that removes the code verification stage, relying solely on CoT analysis. Multi-Dimensional Analysis (MDA) She et al. (2025) is a textual verifier that evaluates each step from five distinct perspectives: restatement, data check, logical consistency, numeric accuracy, and format. Finally, Rephrase–React–Analysis (RRA) Deng et al. (2023) rephrases the current step as a question, solves the question using the ReAct method Yao et al. (2022), and then provides a step-by-step analysis. Implementation details are provided in Appendix A.1.
Training PRMs for TQA.
We use the training sets of WTQ (Pasupat and Liang, 2015) and TabFact (Chen et al., 2020) as our source problem sets for obtaining reasoning paths. For each question table pair, we sample reasoning paths from Qwen2.5-7B-Instruct (Qwen et al., 2025) with a temperature of 0.6. For verification generation with GenPRM, we adopt the default hyper-parameters from Zhao et al. (2025) and perform relative progress estimation (RPE), defined as the ratio between the scores of and . These scores are approximated using Monte Carlo sampling Wang et al. (2024a). We employ QwQ-32B Team (2025) as the rationale generation model to produce both textual analyses and Python code. The normalized probability of the Yes token is taken as the final step reward. If the rewards generated by RPE and the rationale model disagree, the instance is discarded. We also filter out incomplete outputs, instances with improper formatting, and cases where the predicted label is neither Yes nor No. To generate training data for textual verification, QwQ-32B is used. We provide analysis of generated data quality in Appendix A.2. For PRM training, we fully fine-tune DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI et al. (2025), chosen for its large context length, on the filtered dataset. The training data are formatted as standard multi-turn conversations. Training parameters can be found in Appendix A.3.
Baselines and Evaluation.
We compare against the following baselines: Oracle which selects the correct answer from the candidate set, serving as an upper bound on verification performance. Majority@N chooses the most frequent answer among candidates; in the case of a tie, the first prediction is selected. LLM-as-Judge uses the QwQ-32B model as an LLM judge, with the prompt provided in Appendix A.4. We adopt two evaluation paradigms: answer-focused and process-focused. In the answer-focused setting, we apply a best-of- strategy: each sampled reasoning path is scored by verifiers, and the final answer is taken from the path with the highest average step score. Performance is measured using exact match (EM). We report results in both in-domain and out-of-domain scenarios, evaluating on the test sets of WTQ (Pasupat and Liang, 2015), TabFact (Chen et al., 2020), CRT (Zhang et al., 2023), and SciTab (Lu et al., 2023). Dataset statistics are shown in Appendix A.5. In the process-focused settings, we manually annotate 300 instances which are randomly sampled from the four datasets, resulting in 1916 steps (1366 correct and 550 incorrect).
3 Results and Analysis
Figure 2 presents EM of various PRMs and baseline approaches across different path sampling budgets on WTQ and TabFact. Results for step-level verification are shown in Table 1.
GenPRMs achieve promising performance compared to other methods, while falling short of oracle for in-domain settings.
From Figure 2, we observe that GenPRMs consistently delivers the strongest results compared with examined generative PRMs. We attribute it to its integrated code verification component. A detailed analysis is provided in the following paragraph. In comparison to baseline methods, both the 7B and 32B variants of GenPRMs substantially outperform majority voting, underscoring the benefit of process-level verification for answer selection. Nevertheless, the fine-tuned 7B GenPRMs lags behind the 32B LLM-as-a-judge. Though scaling GenPRMs to 32B closes this gap, performance advantages over LLM-as-a-judge remain limited. Across all methods, a substantial performance gap remains compared to oracle, indicating considerable room for improvement in leveraging multiple candidate solutions. We discuss the efficiency of different methods in Appendix A.6.
Incorporating code verifier improves PRM performance in both answer and step correctness.
From the answer-level perspective, a comparison between CoT PRM and GenPRM in Figure 2 shows that GenPRM substantially outperforms CoT PRM. Since CoT PRM is trained on the same data as GenPRM but without the code verification component, this performance gain can be attributed to the incorporation of code verification. From the process-level perspective, we examine step-level correctness by comparing rewards derived from pure textual analysis with those obtained after applying code verification in GenPRM. The results, presented in the last four rows of Table 1, demonstrate that incorporating code verification markedly improves step-level correctness verification across datasets and model sizes.
| Method | WTQ | TabFact | CRT | SciTab |
| LLM-as-a-Judge (32B) | 90.29 | 88.56 | 81.60 | 84.82 |
| Textual PRM (7B) | 81.06 | 80.35 | 71.58 | 74.70 |
| MDA (7B) | 80.79 | 81.25 | 72.10 | 72.61 |
| RRA (7B) | 71.52 | 84.82 | 70.37 | 77.08 |
| GenPRM w/o code (7B) | 80.99 | 80.29 | 71.58 | 75.00 |
| GenPRM (7B) | 91.38 | 90.13 | 77.58 | 84.93 |
| GenPRM w/o code (32B) | 84.79 | 87.50 | 79.80 | 80.92 |
| GenPRM (32B) | 93.71 | 94.72 | 83.22 | 85.71 |
When Code Helps and Hurts.
We conduct a manual analysis to characterize when code-based verification is beneficial and when it is not. Specifically, we review 60 randomly sampled steps on which GenPRM with code and without code disagree (e.g., a step labeled 1 by GenPRM with code but 0 by the model without code). Among these 60 instances, GenPRM with code assigns the correct label in 43 cases and an incorrect label in 17 cases. Within the correctly classified instances, we observe two recurring patterns in which code execution provides substantial benefits: (1) Numerical operations. Executing code supports precise and quantitative reasoning (e.g., counting, aggregation, and threshold checks), thereby reducing ambiguity and limiting error propagation. (2) Targeted retrieval. Programmatic access to specific rows or columns provides accurate relevant evidence and enables verification of individual cell values. We also identify two common failure modes of code-based verification: (1) Exhaustive retrieval. The program retrieves large portions of a table (or even a full table) even when the claim depends on a small subset, implicitly shifting the burden to long-context recall rather than selective filtering. (2) Index misunderstandings. The model sometimes confuses row indexing conventions, inconsistently switching between absolute row indices (including the header) and row indices that exclude the header, which leads to incorrect cell retrieval.
Fine-tuned PRMs with code verification are not robust on out-of-domain datasets.
As shown in Figure 2, GenPRM consistently outperforms majority voting on in-domain datasets. However, this advantage diminishes in out-of-domain settings, with performance dropping to 53 vs.56 on CRT and 60 vs.62 on SciTab, as reported in Table 3. We attribute this gap to differences in problem complexity: CRT contains highly complex questions, while SciTab requires specialized scientific knowledge. These challenging scenarios often lead the policy model to generate reasoning paths containing spurious or noisy steps, which hinder effective verification. An analysis of variance in reasoning complexity across datasets is provided in Appendix A.7. In summary, supervised fine-tuned PRMs are prone to generalization issues. Mitigating this limitation may require more balanced fine-tuning data that reflects varying verification difficulties, or additional post-training techniques to enhance the generalizability of PRMs.
Process-level evaluation highlights the competitiveness of GenPRM.
While the fine-tuned 7B GenPRM under-performs the 32B LLM-as-a-judge in answer evaluation, process-level verification reveals a different trend: 7B GenPRM attains comparable or higher accuracy across most datasets, with the exception of CRT (Table 1). This suggests that strong step-level verification does not necessarily translate into improved answer accuracy. We elaborate on this point in the next paragraph. Moreover, purely text-based PRMs consistently lag behind GenPRM, underscoring the benefit of integrating code verification.
Weak correlation between PRM verification and answer correctness in TQA.
Our manual analysis reveals that average step scores are not strongly correlated with answer correctness. Figure 3 confirms this trend by plotting EM accuracy across bins of average step scores. We observe that solutions with higher average step scores do not necessarily feature higher EM accuracy. A likely reason is that averaging step-level rewards treats all steps as equally informative, regardless of their causal contribution to the final answer. Consequently, (i) chains containing many irrelevant but locally correct steps receive inflated scores, whereas (ii) chains with a local error that does not affect the final answer are over-penalized. Examples of both cases are provided in Appendix A.8. Unlike in math, reasoning chains in TQA are more loosely linked, so an error at step does not necessarily propagate to subsequent steps or the final prediction. This reduces the effectiveness of step verification for answer selection. Nevertheless, PRMs remain valuable; the core bottleneck lies in the current answer-focused evaluation paradigm and the absence of process-annotated evaluation datasets for TQA. We advocate future work on process-level evaluation and modeling approaches that strengthen step dependencies and assess each step by both its local correctness and its causal contribution to the final answer, for example, by filtering irrelevant steps.
Potential Alternative Aggregation Strategies.
We observe that simple uniform averaging is suboptimal, as it implicitly assumes that all reasoning steps contribute equally to the final answer; this assumption allows irrelevant or non-causal steps to influence the estimated solution quality and obscure the impact of steps that are truly determinative. Beyond uniform averaging, we explore alternative aggregation strategies that more accurately reflect answer quality in TQA. A simple and intuitive approach is to assign greater weight to the final reasoning steps. When evaluated using a 32B coding verifier, this strategy yields consistent improvements of +3.56% EM on WTQ, +1.14% on TabFact, +4.55% EM on CRT, and +2.69% on SciTab. In addition, relevance- and causality-aware aggregation strategies represent promising directions for future study. A relevance-aware strategy assigns a relevance score to each step, indicating the step’s relevance to solving the problem. Similarly, a causality-aware strategy assigns a causality score to each step, indicating whether the step influences the final outcome. Incorporating these weights when computing a solution score can down-weight steps that do not contribute to the final answer, thereby leading to a score that more faithfully reflects answer quality.
4 Conclusions
In this paper, we examined four generative PRMs on TQA and compared them to strong baselines such as LLM-as-a-judge and majority vote from both answer and process evaluation perspectives. We found that PRMs that combine textual and code verification can aid solution selection but struggle to generalize to out-of-domain data. Our analysis suggests that the primary bottleneck is not step-level verification, but rather the aggregation of step rewards. Uniform averaging is inadequate for TQA as it assigns equal weight to all steps, regardless of their relevance to the problem or their causal influence on the final answer. Future work can therefore investigate relevance-aware and causality-aware aggregation strategies to generate solution scores that more faithfully reflect answer quality.
Limitations
To examine the effectiveness of generative PRMs over semi-structured data, we choose table question answering as a proxy task for initial exploration. Nevertheless, there exist other tasks involving semi-structured data, such as table generation or table summarization. Future studies can explore these tasks to validate the effectiveness of PRMs. Moreover, we mainly train PRM using the Qwen models, following the practice of PRM training in previous work Zhao et al. (2025). Though we experiment with varying model sizes, results from a different model family might be insightful.
Appendix A Appendix
A.1 Implementation and Adaptation Details
We adapt five perspectives from She et al., 2025—previous steps analysis, now step analysis, data source analysis, consistency analysis, and calculation analysis—into ours: restatement, data check, logical consistency, numeric accuracy, and format, as shown in Figure 10. For RRA, we follow Deng et al., 2023: the PRM first rephrases the current step, then solves the rephrased question using the ReAct paradigm Yao et al. (2022), and finally provides a step-by-step analysis conditioned on the current step, the rephrased problem, and the ReAct trace, as shown in Figure 11.
A.2 Data Quality Analysis
| Label | After Filtering | Before Filtering | ||
| #Steps | % | #Steps | % | |
| Correct | 254 | 94.42 | 240 | 84.21 |
| Incorrect | 15 | 5.58 | 45 | 15.79 |
Training data quality may partially account for the limited performance of GenPRM on TQA. To examine this, we randomly sample 50 reasoning paths before and after the GenPRM filtering procedure and manually assess their correctness. As shown in Table 2, the filtering process effectively removes incorrect steps. After filtering, which produces the data used for PRM training, the step-level accuracy reaches 94%, indicating overall high data quality. This suggests that data noise is unlikely to be the primary bottleneck for GenPRM performance on TQA.
A.3 Training Parameters
Models were fine-tuned with full-parameter. Training proceeded for a single epoch with a batch size of 64 and a maximum sequence length of 2,048 tokens. We initialized the learning rate at and applied a cosine decay schedule. Optimization used AdamW with bfloat16 (bf16) precision.
A.4 Prompts
Prompts for generating reasoning paths for WTQ, TabFact, CRT and SciTab are shown in Figure 6, 7, 6 and 8, respectively. Prompts for generating rationales for a textual PRM are presented in Figure 9. Prompts for generating rationales for Multi-Dimensional Analysis and Rephrase, React, Analysis are shown in Figure 10 and Figure 11, respectively. Refer to Figure 12 for the prompt for data generation for GenPRM. Lastly, Figure 13 shows the prompt for collecting data for LLM-as-a-judge.
| Method | WTQ | TabFact | CRT | SciTab |
| Upper Bound | 82.52 | 97.72 | 71.01 | 88.39 |
| Pass@1 | 57.65 | 85.24 | 51.09 | 57.10 |
| Majority@8 | 64.86 | 88.52 | 56.45 | 62.58 |
| LLM-as-a-Judge (7B) | 59.06 | 85.51 | 50.13 | 58.00 |
| LLM-as-a-Judge (32B) | 67.94 | 91.48 | 55.49 | 64.62 |
| CoT-style PRM (7B) | 62.88 | 88.44 | 51.51 | 60.07 |
| CoT-style PRM (32B) | 67.48 | 91.86 | 55.08 | 63.36 |
| Multi-dimensional (7B) | 62.18 | 87.89 | 51.94 | 59.96 |
| Rephrase-React-Analysis (7B) | 62.66 | 88.69 | 52.74 | 59.93 |
| GenPRM (7B) | 66.09 | 90.20 | 53.31 | 60.64 |
| GenPRM (32B) | 68.92 | 92.42 | 56.00 | 63.78 |
A.5 Dataset Statistics and Licenses
The statistics of datasets we used are shown in the Table 4.
| Datasets | #instances | Domain |
| WTQ | 4344 | Wikipedia |
| TabFact | 2024 | Wikipedia |
| SCITAB | 1224 | scientific paper |
| CRT | 728 | Wikipedia |
WTQ Pasupat and Liang (2015) is under the license of CC-BY-SA-4.0222https://creativecommons.org/licenses/by-sa/4.0/. TabFact Chen et al. (2020), CRT Zhang et al. (2023) and SCITAB Lu et al. (2023) are under the MIT333https://opensource.org/license/mit license.
A.6 Time Efficiency.
| WTQ | TabFact | CRT | SciTab | |
| Textual PRM | 61.56 | 80.60 | 56.82 | 90.32 |
| Multi-Dimensional | 82.75 | 88.86 | 83.32 | 112.64 |
| Rephrase-React-Analysis | 150.78 | 133.81 | 208.07 | 163.73 |
| GenPRM w/o code (7B) | 107.22 | 122.41 | 154.84 | 188.31 |
| GenPRM (7B) | 222.93 | 228.38 | 290.30 | 469.45 |
| GenPRM w/o code (32B) | 185.96 | 159.95 | 255.65 | 191.29 |
| GenPRM (32B) | 508.66 | 433.21 | 608.91 | 529.66 |
Despite dataset-specific variation, PRMs without code (Textual, MDA, RRA, and GenPRM without execution) are consistently cheaper than GenPRM. Code execution introduces significant runtime overhead, especially with larger models. This suggests that integrating code improves performance at the cost of higher latency and longer inference time.
A.7 Variance in Reasoning Complexity Across Datasets
Table 6 reveals a clear in-/out-of-domain split in path structure. WTQ and TabFact predominantly exhibit consistent trajectories (86% and 81%), i.e., monotone label sequences with at most one change-point. By contrast, CRT and SciTab show far more inconsistent paths (37% and 50%), with multiple flips between correct and incorrect steps. This suggests fewer spurious/noisy steps in the in-domain sets (WTQ/TabFact) and a higher noise ratio in CRT/SciTab.
A.8 Examples of Inflated Scores and Over-Penalization
| Reasoning Paths | WTQ | TabFact | CRT | SciTab |
| Total | 100 | 100 | 100 | 100 |
| Consistent | 86 | 81 | 63 | 50 |
| Inconsistent | 14 | 19 | 37 | 50 |